The income from operations and the amount of invested assets in each division of Beck Industries are as follows: I...

  1. The income from operations and the amount of invested assets in each division of Beck Industries are as follows:

    Income from Operations Invested Assets
    Retail Division $48,600 $270,000
    Commercial Division 52,500 250,000
    Internet Division 169,400 770,000

    Assume that management has established a 10% minimum acceptable return for invested assets.

    a. Determine the residual income for each division.

    Retail Division Commercial Division Internet Division
    Income from operations $48,600 $52,500 $169,400
    Minimum acceptable income from operations as a percent of invested assets
    Residual income $ $ $

    b. Which division has the most residual income
Answer #1
Retail Commercial Internet
Income 48,600 52,500 169,400
Min acceptable income (270,000*10%) = 27,000 (250,000*10%) = 25,000 (770,000*10%) = 77,000
Residual income 21,600 27,500 92,400

Internet division has most residual income